11. R.A Mona receives and produces an order. What is the companys cycle efficiency assuming the following times were measured during production of this order Process time: a) 10% b) 28% c) 42% d) 20% e) 50% .7 days Inspection time: .25 days Move time: 1.05 days Wait time: 5 days
Answer #1
Manufacturing cycle efficiency=Process time/Manufacturing cycle time
Process time=0.7 days
Manufacturing cycle time=Process time+Inspection time+Move time+Wait time=0.7+0.25+1.05+0.5=2.5 days
Manufacturing cycle efficiency=0.7/2.5=0.28=28%
Answer is b) 28%
I appreciate your ratings